A school has 731 students. Prove that there must be at least three students that have the same birthday assuming that no one has a birthday on Feb 29.
Can anybody please help us with this question?
We would be very grateful
Thank you💜

Found 2 solutions by greenestamps, ikleyn:
Answer by greenestamps(13200)

Look at the question this way: you are trying to find as many people as you can without any three of them having the same birthday.

Assuming nobody has a birthday on February 29, you have 365 days in the year that can be a person's birthday.

The worst case scenario is that you find exactly 2 people who have birthdays on each of the 365 days of the year. You still have no 3 people with the same birthday; and the number of people you have will be 2*365 = 730.

But then the 731st person has to have a birthday on one of those 365 days; that means you will have three people with the same birthday.

And of course it is unlikely that you will easily find 730 people without having 3 of them with the same birthday; that means with 731 people you would almost certainly have more than one date that is a birthday for 3 people; and you could easily have one birthday that is shared by MORE THAN 3 of the people.

Answer by ikleyn(52786)

.

There is so called  "the pigeonhole principle"  in Math:


    If 7 pigeons are placed in 6 holes, then at least one hole contains 2 or more pigeons. 


In this joking form it is obvious and does not require more detailed proofs / explanations.




Further, in more general form, 


    if (n+1) pigeons are placed in "n" holes, then there is at least one hole containing 2 or more pigeons. 


In this form it is obvious, again, and does not require more detailed proofs / explanations.




Now let me formulate even more general THEOREM.


    If there are  n*m+1  items in  "m" containers, then there is at least one container containing  "n+1"  or more  items.



The proof is in three lines.


    If there is NO such a container, then the total number of items in "m" containers is NOT MORE than  n*m.


    It  CONTRADICTS  to the given part (!)


    This contradiction proves the statement.




Now let's return to our problem.


We have 731 = 2*365 + 1  students  (= items),  and

        365 days in the year (that are "containers" in this case).


From the Theorem, there is at least one container containing (2+1) = 3 or more items.

In other words, there is at least one day (one date in an year), when 3 or more students celebrate their birthdays.


Solved.


-----------------

In  MATHEMATICS  (notice all letters are capital,  which means that I am talking about  TRUE  Math)

this principle is called  "the Dirichlet's principle",  after the famous mathematician  Johann  Peter  Gustav  Lejeune  Dirichlet
(1805 - 1859).


About  Dirichlet,  see these Internet articles
http://www-history.mcs.st-and.ac.uk/Biographies/Dirichlet.html
https://en.wikipedia.org/wiki/Peter_Gustav_Lejeune_Dirichlet


About  "the Dirichlet's principle",  see this Internet article
https://en.wikipedia.org/wiki/Pigeonhole_principle